OpenStudy (anonymous):

A total electric charge of 4.60 nC is distributed uniformly over the surface of a metal sphere with a radius of 22.0 cm. The potential is zero at a point at infinity. Find the value of the potential at 56.0 cm from the center of the sphere.

OpenStudy (anonymous):

use the eqaution of electric potential v\[v=kq \div r \ where k=\[9\times10^{9}\]

OpenStudy (anonymous):

\[\Delta V = Vb - Va = -\int\limits_{b}^{a} E ds\] \[E = kQ ^ r/r ^{2} ds\] \[-\int\limits_{a}^{b} kQ ^r/r ^{2} ds \rightarrow -kQ \int\limits_{a}^{b} dr/ r ^{2}\] so the equation become : \[kQ ( 1/rb -1/ra)\] when : \[r =\infty\] so the V=0 the final eq : \[V potential charge (r) = kQ/r\]

OpenStudy (anonymous):

agreed